OpenStudy (anonymous):

If f(x) = x + 3, find f(-8). A) 11 B) -5 C) -11 D) 5

OpenStudy (anonymous):

11

OpenStudy (anonymous):

f(8) = 8 + 3 8 + 3 = A) 11

jimthompson5910 (jim_thompson5910):

f(x) = x + 3 f(-8) = -8 + 3 f(-8) = -5

OpenStudy (anonymous):

oh it was a -8 my bad lol

OpenStudy (anonymous):

o.o

jimthompson5910 (jim_thompson5910):

tsitrond lost a sign, but it happens to all of us

OpenStudy (anonymous):

jiim is right

OpenStudy (anonymous):

it's -5

OpenStudy (anonymous):

Thanks :D